OpenStudy (anonymous):

In an election for school council, there were 234 yes and 586 no votes were cast. Write the fraction of no votes

OpenStudy (cwrw238):

total votes = 234 + 586 = 820 fraction of no votes = 586 / 820

OpenStudy (cwrw238):

can you simplify that?

OpenStudy (cwrw238):

you can divide both top and bottom by 2 = 293 / 410

OpenStudy (anonymous):

Thank you

OpenStudy (cwrw238):

yw